Find the average value of f(x,y)=5 x^2 y^4 over the rectangle R with vertices (−2,0),(−2,6),(2,0),(2,6).
Average value =
1
Expert's answer
2020-10-25T19:18:21-0400

The rectangle "R" with vertices "(-2,0),(-2,6),(2,0)" and "(2,6)" is as shown in the figure below:





The area of the rectangle is,


"A(R)=(2+2)(6-0)=(4)(6)=24"


The average value of function over the rectangle "R" is evaluated as,


"f_{ave}=\frac{1}{A(R)}\iint_{R}f(x,y)dA"


"=\frac{1}{24}\int_{-2}^{2}\int_{0}^{6}5x^2y^4dydx"


"=\frac{5}{24}[\frac{x^3}{3}]_{-2}^{2}[\frac{y^5}{5}]_{0}^{6}"


"=\frac{5}{24}(\frac{8}{3}+\frac{8}{3})(\frac{6^5}{5})"


"=\frac{5}{24}(\frac{16}{3})(\frac{7776}{5})"


"=1728"


Therefore, the average value of the function over the rectangle is "f_{ave}=1728" .